OpenStudy (anonymous):

Find the value of x so that the line passing through (-2, 7) and (x, 2) has a slope of -1/3

OpenStudy (anonymous):

(2-7)/(x+2) = -1/3 solve for x

OpenStudy (anonymous):

if you prefer -5/(x+2) = -1/3 making 15 = x + 2

OpenStudy (ness9630):

Game and set^

OpenStudy (anonymous):

so the answer is 13=x correct @satellite73